OpenStudy (anonymous):

Could someone please explain how to find the indefinite integral of csc(x) ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

\(\csc x=\frac{\csc x(-\csc x+\cot x)}{(-\csc x+\cot x)}\) Now use substitution \(u=-\csc x+\cot x\). Then the integral become \(\int du/u=\ln |u|+C=\ln|-\csc x+\cot x|+C\)

OpenStudy (anonymous):

so you have to multiply by a factor of the derivative of csc(x) divided by itself, or 1, to solve?

OpenStudy (watchmath):

Well, first yes we multiply by 1 so it won't change anything. But \(-\csc x+\cot x\) is not the derivative of \(\csc x\). What I can say that we multiply that weird expression just to make it work. How the first person come up with this trick I have no idea :D.

OpenStudy (anonymous):

oops i did not take notice of the plus sign between them, now i see, thanks for the explanation